# Universal bounds on optimization of free energy harvesting

### 20 Dec 2023 v3

@@ Abstract
Harvesting free energy from the environment is essential for the operation of many biological and artificial systems. We investigate the maximum rate of harvesting achievable by optimizing a set of reactions in a Markovian system, possibly given topological, kinetic, and thermodynamic constraints. We show that the maximum harvesting rate can be expressed as a variational principle, which we solve in closed-form for three physically meaningful regimes. Our approach is relevant for optimal design and for quantifying efficiency of existing reactions. Our results are illustrated on bacteriorhodopsin, a light-driven proton pump from Archae, which is found to be close to optimal under realistic conditions.
@@
Many molecular systems, both biological and artificial, harvest free energy from their environments. Biological organisms require free energy to grow and replicate [1, 2], and many species undergo selection for increased energy harvesting [3, 4, 5, 6]. Artificial harvesting systems have also been constructed and optimized in the field of synthetic biology [7, 8, 9, 10, 11, 12, 13, 14]. The optimization of free energy harvesting is thus a central problem both in biology and engineering.
@@
As an example, consider a harvesting system such as a biological metabolic network that converts glucose to ATP [15]. Suppose that the kinetic and thermodynamic parameters of one or more reactions can be optimized, either by natural selection or artificial design. What is the maximum rate of free energy harvesting that the network can achieve, and what are the kinetic and thermodynamic parameters that achieve it? These questions are relevant both for optimization and for quantifying the efficiency of existing systems.
In this Letter, we use techniques from stochastic thermodynamics to derive bounds on maximum rate of free energy harvesting achievable in steady state. We also identify optimal steady states and kinetic parameters that achieve those bounds. Our approach is universal in that it is not restricted to any particular model. At the same time, it allows for the incorporation of domain-specific constraints on network topology, kinetic timescales, and thermodynamics.
@@
Our approach is well suited for the kinds of small-scale systems usually considered in stochastic thermodynamics [16], where assumptions of local detailed balance and steady state are justified. The steady-state assumption is reasonable in many molecular systems, where there is a separation of timescales between internal relaxation and environmental change.
We illustrate our results on bacteriorhodopsin [17]. Bacteriorhodopsin is an important photosynthetic system found in Archae and having relatives in bacteria [18, 19], and it is also used in many artificial light-harvesting systems [7, 9, 8, 11]. We demonstrate that, under normal operating conditions, the bacteriorhodopsin system is remarkably efficient.
Our main result is formulated as a variational principle which must be solved numerically in the general case. In the second part of this Letter, we derive closed-form solutions for three physically meaningful regimes. These solutions illustrate how optimal harvesting reflects the “alignment” between free energy input and relaxation dynamics. We illustrate them on an idealized model of a unicyclic system.
Setup.— We consider a system with $n$ mesostates described by the distribution $\bm{p}=(p_{1},\dots,p_{n})$. The distribution evolves according to the master equation $\dot{p}_{i}=\sum_{j}R_{ji}p_{j}$, where $R_{ji}$ is the transition rate $j\to i$ ($R_{ii}=-\sum_{j}R_{ji}$). Usually $\bm{p}$ represents a probability distribution of a stochastic system with Markovian dynamics [20, 21]. However, it may also represent chemical concentrations in a deterministic chemical reaction network with pseudo-first-order reactions (e.g., an enzymatic cycle [22, 23]).
The system is coupled to a heat bath at inverse temperature $\beta=1/k_{B}T$, an internal free energy reservoir, and a nonequilibrium environment that serves as an external source of free energy. For example, in a metabolic network, one may consider an internal reservoir of ATP and an external source of glucose. The system has nonequilibrium free energy $\mathcal{F}(\bm{p})=\sum_{i}p_{i}f_{i}-{\beta}^{-1}S(\bm{p})$, where $S(\bm{p}):=-\sum_{i}p_{i}\ln p_{i}$ is the Shannon entropy and $f_{i}$ is the internal free energy of mesostate $i$ [24].
We focus on the behavior of the system in the steady-state distribution $\bm{\pi}$, which satisfies $R\bm{\pi}=0$. Our aim is to maximize the harvesting rate at which free energy flows into the internal reservoir in steady state. To investigate this question, we separate the dynamics of the system into two kinds of processes, termed baseline and control. Control processes are those that are optimized to increase harvesting, while the baseline processes are held fixed.
@@
The particular separation of baseline/control generally depends on domain knowledge about the system and the scientific question. For example, to study the efficiency of a particular reaction in a metabolic network, that reaction may be treated as control while the other reactions are baseline. The baseline/control separation is similar to the distinction in control theory between “plant” (a given system with fixed dynamical laws) and “controller” (the part that undergoes optimization) [25]. In stochastic thermodynamics, the baseline/control distinction has been used to study autonomous Maxwellian demons [26, 27], counterdiabatic driving [28], and the cost of maintaining a nonequilibrium steady state [29, 30].
@@
We make one important assumption in our analysis: the control processes are only coupled to the heat bath and internal free energy reservoir, but not directly to the external source of free energy. This means that control can only increase harvesting by interacting with the baseline, rather than directly increasing the inflow of free energy from the external source. For example, in a metabolic network, control processes cannot directly increase the import of glucose, but they can affect the rate at which glucose is converted into ATP by optimizing intermediate reactions and mechanisms. Control processes may be driven by the internal reservoir (e.g., coupled to ATP hydrolysis) or undriven (e.g., enzymes).
@@
The free energy increase due to control, ˙$\dot{\mathcal{F}}^{c}(\bm{p})$, is defined analogously but using rate matrix $R^{c}$. The harvesting rate is the rate at which free energy flows to the internal reservoir. The harvesting rate due to baseline and control is
$\dot{\mathcal{G}}^{b}(\bm{p})=\sum_{i,j}p_{i}R^{b}_{ji}{g}^{b}_{ji}+\sum_{i}p_{i}\dot{g}^{b}_{j}\;,\quad\dot{\mathcal{G}}^{c}(\bm{p})=\sum_{i,j}p_{i}R^{c}_{ji}{g}^{c}_{ji}\,.$
Here ${g}^{b}_{ji}$ is the free energy increase in the internal reservoir due to a single baseline transition $i\to j$ and $\dot{g}^{b}_{i}$ is the rate of free energy flow to the internal reservoir due to internal transitions within $i$ (assuming $i$ is a coarse-grained mesostate). Similarly, ${g}^{c}_{ji}$ is the free energy increase in the internal reservoir due to control transition $i\to j$. For simplicity, we assume that control cannot exchange free energy with the internal reservoir due to internal transitions within $i$. Negative values of ${g}^{b}_{ji},\dot{g}^{b}_{i},{g}^{c}_{ji}$ indicate driving done on the system by the internal reservoir.
For a concrete example of how $(R^{b},R^{c},\bm{f},\bm{g}^{b})$ are defined for a real biomolecular system, see the bacteriorhodopsin example below and SM2 [31].
@@
Maximum harvesting rate.— Suppose that the combined rate matrix $R=R^{b}+R^{c}$ has the steady-state distribution $\bm{\pi}$, which satisfies $R^{b}\bm{\pi}+R^{c}\bm{\pi}=0$. The total steady-state harvesting rate due to baseline and control is
@@
We seek to maximize this harvesting rate by varying the parameters of the control processes $(R^{c},\bm{g}^{c})$ while holding the baseline parameters ($\bm{f},R^{b},\bm{g}^{b},\dot{\bm{g}}^{b}$) fixed.
To derive a bound on (3), we re-express it as
@@
Here $\dot{\Sigma}({\boldsymbol{J}^{c}})=\sum_{i\neq j}{J^{c}_{ji}}\ln({J^{c}_{ji}}/{J^{c}_{ij}})\geq 0$ is the Schnackenberg formula for the entropy production rate (EPR) [20], where ${J^{c}_{ji}}=\pi_{i}R^{c}_{ji}\geq 0$ is the one-way probability flux due to control transition $i\to j$. See SM1.1 [31] for details of this derivation, which makes use of the steady-state condition and LDB (2).
We then maximize (4) with respect to the steady-state distribution $\bm{\pi}$ and control fluxes ${\boldsymbol{J}^{c}}$. Note that varying $\bm{\pi}$ and ${\boldsymbol{J}^{c}}$ is equivalent to varying the control rate matrix via $R^{c}_{ji}={J^{c}_{ji}}/\pi_{i}$ and control driving ${g}^{c}_{ji}$ via (2). While maximizing, we also ensure that $\bm{\pi}$ satisfies the steady-state condition $R^{b}\bm{\pi}+R^{c}\bm{\pi}=0$, which can be expressed as the linear constraint $R^{b}\bm{\pi}+\mathbb{B}{\boldsymbol{J}^{c}}=0$. Here ${\boldsymbol{J}^{c}}$ is treated as a vector in $\mathbb{R}^{n^{2}}$ and $\mathbb{B}\in\mathbb{R}^{n\times n^{2}}$ is the incidence matrix with entries $\mathbb{B}_{k,ij}=\delta_{ki}-\delta_{kj}$, which guarantees $R^{c}\bm{\pi}=\mathbb{B}{\boldsymbol{J}^{c}}$.
@@
$\mathscr{G}$ $=\sup_{(\bm{p},{\boldsymbol{J}})\in\Lambda:\mathbb{B}{\boldsymbol{J}}=-R^{b}\bm{p}}\,\dot{\mathcal{F}}^{b}(\bm{p})+\dot{\mathcal{G}}^{b}(\bm{p})-\dot{\Sigma}({\boldsymbol{J}})/\beta.$ (5)
In this expression, $\Lambda$ is the feasible set of distributions $\bm{p}$ and control fluxes ${\boldsymbol{J}}$. At a minimum, $\Lambda$ ensures the validity of the distribution $\bm{p}$ and the fluxes ${\boldsymbol{J}}$ via the linear constraints $\sum p_{i}=1$, $p_{i}\geq 0$, and $J_{ji}\geq 0$. We write $\sup$ instead of $\max$ because the set of allowed fluxes is potentially unbounded.
Importantly, the feasible set $\Lambda$ can include additional convex constraints, which may introduce topological, kinetic, thermodynamic, etc. restrictions on the control processes. Topological constraints restrict which transitions can be controlled; e.g., $J_{ji}=0$ ensures that control does not use transition $i\to j$). Kinetic constraints restrict timescales of control processes, as might reflect underlying physical processes like diffusion; e.g., an upper bound on control transition rate $R^{c}_{ji}=J_{ji}/p_{i}\leq\kappa$ can be enforced via $J_{ji}\leq p_{i}\kappa$. Thermodynamic constraints bound the affinity [20] of control transitions; e.g., $J_{ji}e^{-\mathcal{A}}\leq J_{ij}\leq J_{ji}e^{\mathcal{A}}$ ensures that $|\ln(J_{ij}/J_{ji})|\leq\mathcal{A}$. The above examples all involve linear constraints. An example of a nonlinear, but still convex, constraint is an upper bound on the EPR incurred by control, $\dot{\Sigma}({\boldsymbol{J}^{c}})\leq\dot{\Sigma}^{c}_{\max}$.
Eq. (5) is our first main result. Importantly, $\mathscr{G}$ is defined purely in terms of the thermodynamic and kinetic properties of the baseline process, along with desired constraints encoded in $\Lambda$. Thus, $\mathscr{G}$ is the maximum steady-state harvesting rate that can be achieved by any allowed control processes, given a fixed baseline.
Eq. (5) involves the maximization of a concave objective (equivalently, minimization of a convex objective) given convex constraints (SM1.2 [31]). This is a convex optimization problem that can be efficiently solved using standard numerical techniques [34]. The optimization also returns an optimal steady-state distribution $\bm{p}^{*}$ and control fluxes ${{\boldsymbol{J}}}^{*}$ that achieve the maximum harvesting rate $\mathscr{G}$ (or come arbitrarily close to achieving it). These fix the optimal control rate matrix via $R^{c*}_{ji}={J_{ij}^{*}}/\pi_{i}$. Thus, our optimization specifies an upper bound on harvesting as well as the optimal control strategy that achieves this bound.
@@
The optimal $\bm{p}^{*}$ is unique as long as the baseline rate matrix is irreducible. The optimal control rate matrix is very fast ($R^{c*}\to\infty$) and obeys detailed balance for $\bm{p}^{*}$, $R^{c*}_{ji}{p}^{*}_{i}=R^{c*}_{ij}{p}^{*}_{j}$ (SM1.3 and SM1.4 [31]).
Figure 1: Top left: Bacteriorhodopsin is a biomolecular light-powered proton pump [17]. Right: comparison of the actual harvesting rate $\dot{\mathcal{G}}^{\mathrm{tot}}$ at different electrical potentials, versus maximum rate $\mathscr{G}$ achieved by optimizing five intermediate transitions (colors in top left). Inset: transition efficiencies $\dot{\mathcal{G}}^{\mathrm{tot}}/\mathscr{G}$. Lower left: cycle steady state $\bm{\pi}$ and optimal distribution $\bm{p}^{*}$ for $N\leftrightarrow O$ as control ($\Delta\psi=120$ mV).
Bacteriorhodopsin.— We illustrate our results using bacteriorhodopsin, a light-driven proton pump from Archae [17].
We define a thermodynamically consistent model of the bacteriorhodopsin cycle using published thermodynamic [35] and kinetic [36] data (see SM2 [31]). The system operates in a cyclical manner, absorbing a photon and pumping a proton during each turn of the cycle (Fig. 1, top left). Specifically, the transition $M_{1}\to M_{2}$ pumps a proton out of the cell. This stores free energy in the internal reservoir (the membrane electrochemical potential),

### 12 Mar 2024 v4

# Optimization of nonequilibrium free energy harvesting illustrated on bacteriorhodopsin

@@ Abstract
Harvesting free energy from the environment is essential for the operation of many biological and artificial systems. We use techniques from stochastic thermodynamics to investigate the maximum rate of harvesting achievable by optimizing a set of reactions in a Markovian system, possibly under various kinds of topological, kinetic, and thermodynamic constraints. This question is relevant for the optimal design of new harvesting devices as well as for quantifying the efficiency of existing systems. We first demonstrate that the maximum harvesting rate can be expressed as a constrained convex optimization problem. We illustrate it on bacteriorhodopsin, a light-driven proton pump from Archaea, which we find is close to optimal under realistic conditions. In our second result, we solve the optimization problem in closed-form in three physically meaningful limiting regimes. These closed-form solutions are illustrated on two idealized models of unicyclic harvesting systems.
I Introduction
@@
Many molecular systems, both biological and artificial, harvest free energy from their environments. Biological organisms require free energy to grow and replicate [1, 2], and many species undergo selection for increased harvesting [3, 4, 5, 6]. Artificial harvesting systems have also been constructed and optimized in the field of synthetic biology [7, 8, 9, 10, 11, 12, 13, 14]. The optimization of free energy harvesting is thus a central problem both in biology and engineering.
@@
As an example, consider a harvesting system such as a biological metabolic network that converts glucose to ATP [15]. Suppose that the kinetic and thermodynamic parameters of one or more reactions can be optimized, either by natural selection or artificial design. What is the maximum rate of free energy harvesting that the network can achieve, and what are the kinetic and thermodynamic parameters that achieve it? These questions are relevant both for design of optimal harvesting devices and for quantifying the efficiency of existing systems.
@@
In this paper, we use techniques from stochastic thermodynamics to derive bounds on maximum rate of free energy harvesting. We consider a harvesting system in nonequilibrium steady state which is coupled to an external source of free energy, an internal free energy reservoir, and a heat bath. The setup is well-suited for studying the kinds of small-scale systems usually considered in stochastic thermodynamics [16], where assumptions of local detailed balance and steady state are justified. The steady-state assumption is reasonable in many molecular systems, where there is a separation of timescales between internal relaxation and environmental change.
@@
We suppose that the system’s dynamics can be separate into two kinds of processes, termed baseline and control. The baseline processes, which are held fixed, mediate the coupling to the external source of free energy. Control refers to all other processes which can be optimized for increasing the harvesting rate at which free energy flows into the internal reservoir. The particular separation of baseline/control generally depends on domain knowledge about the system and the scientific question at hand. For example, to study the efficiency of a particular reaction in a metabolic network, that reaction may be treated as control while the other reactions are baseline. The baseline/control separation is similar to the distinction in control theory between “plant” (a given system with fixed dynamical laws) and “controller” (the part that undergoes optimization) [17].
In our first set of results, we demonstrate that the optimization of the harvesting rate can be expressed as a convex optimization problem. The optimal solution of this problem determines both the maximum harvesting rate and the specific control processes that achieve that maximum. Importantly, the optimization can also account for various types of constraints on the possible network topology, kinetic timescales, and thermodynamic forces of the control processes.
Figure 1: Left: Bacteriorhodopsin is a biomolecular free energy harvesting machine [18], illustrated in its trimer configuration by D. Goodsell (CC-BY-4.0) [19]. Right: during each turn of the bacteriorhodopsin photocycle, the molecule absorbs a photon and pumps a proton against the cell’s membrane potential.
We illustrate our results on bacteriorhodopsin (Fig. 1), a proton-pumping membrane protein. Bacteriorhodopsin is a photosynthetic system found in Archaea, with close relatives in bacteria [20, 21]. It is also used in many artificial light-harvesting systems [7, 9, 8, 11]. Using published thermodynamic and kinetic data, we develop a thermodynamically consistent stochastic model of bacteriorhodopsin. We demonstrate that, under normal operating conditions, the bacteriorhodopsin system is remarkably efficient.
Our main result is formulated as a convex optimization problem which must be solved numerically in the general case. In the second part of this paper, we derive closed-form solutions of this problem for three physically meaningful regimes: the weakly-driven linear response regime, the irreversible deterministic regime, and the intermediate near-deterministic regime. These solutions illustrate how optimal harvesting reflects the “alignment” between free energy input and relaxation dynamics. We illustrate these closed-form solutions on two unicyclic systems, which may be interpreted as idealized models of two types of nonequilibrium harvesting devices.
We finish our paper with a brief Discussion. There we relate our approach to previous work, including maximization of power output in steady-state engines and flux balance analysis. We also propose directions for future research.
II Setup
We consider a system with $n$ mesostates described by the distribution $\bm{p}=(p_{1},\dots,p_{n})\in\mathbb{R}_{+}^{n}$. The distribution evolves according to the master equation $\dot{p}_{i}=\sum_{j}R_{ij}p_{j}$, where $R_{ji}$ is the transition rate $i\to j$ ($R_{ii}=-\sum_{j}R_{ji}$). Usually $\bm{p}$ represents a probability distribution of a stochastic system with Markovian dynamics [22, 23]. However, under an appropriate choice of units, it may also represent chemical concentrations in a deterministic chemical reaction network with pseudo-first-order reactions, such as an enzymatic cycle [24, 25].
The system is coupled to a heat bath at inverse temperature $\beta=1/k_{B}T$, an internal free energy reservoir, and a nonequilibrium environment that serves as an external source of free energy. For example, in a metabolic network, one may consider an internal reservoir of ATP and an external source of glucose. The system has nonequilibrium free energy
$\mathcal{F}(\bm{p})=\sum_{i}p_{i}f_{i}-{\beta}^{-1}S(\bm{p})\,,$ (1)
where $S(\bm{p}):=-\sum_{i}p_{i}\ln p_{i}$ is the Shannon entropy and $f_{i}$ is the internal free energy of mesostate $i$ [26].
@@ II Setup
As mentioned in the Introduction, we suppose that the dynamics of the system can be separated into baseline and control processes. We make one important assumption in our analysis: the control processes are only coupled to the heat bath and internal free energy reservoir, but not directly to the external source of free energy. This means that control can only increase harvesting by interacting with the baseline, rather than directly increasing the inflow of free energy from the external source. For example, in a metabolic network, control processes cannot directly increase the import of glucose, but they can affect the rate at which glucose is converted into ATP by optimizing intermediate reactions and mechanisms. Control processes may be driven by the internal reservoir (e.g., coupled to ATP hydrolysis) or undriven (e.g., enzymes).
@@
The increase due to control processes is defined analogously but using rate matrix $R^{c}$,
˙$\dot{\mathcal{F}}^{c}(\bm{p})=\sum_{i,j}p_{i}R^{c}_{ji}(f_{j}+{\beta}^{-1}\ln p_{j})\,.$ (3)
The harvesting rate is the rate at which free energy flows to the internal reservoir. The harvesting rate due to baseline processes is
$\dot{\mathcal{G}}^{b}(\bm{p})$ $=\sum_{i,j}p_{i}R^{b}_{ji}{g}^{b}_{ji}+\sum_{i}p_{i}\dot{g}^{b}_{i}\,,$ (4)
where ${g}^{b}_{ji}$ is the free energy increase in the internal reservoir due to a single baseline transition $i\to j$ and $\dot{g}^{b}_{i}$ is the rate of free energy flow to the internal reservoir due to internal transitions within $i$ (assuming $i$ is a coarse-grained mesostate). Similarly, the harvesting rate due to control processes is
$\dot{\mathcal{G}}^{c}(\bm{p})$ $=\sum_{i,j}p_{i}R^{c}_{ji}{g}^{c}_{ji}\,,$ (5)
where ${g}^{c}_{ji}$ is the free energy increase in the internal reservoir due to control transition $i\to j$. For simplicity, we assume that control cannot exchange free energy with the internal reservoir due to internal transitions within $i$. Negative values of ${g}^{b}_{ji},\dot{g}^{b}_{i},{g}^{c}_{ji}$ indicate driving done on the system by the internal reservoir.
For a concrete example of how $(R^{b},R^{c},\bm{f},\bm{g}^{b},\dot{\bm{g}}^{b},\bm{g}^{c})$ are defined for a real biomolecular system, see the bacteriorhodopsin example below and SM-II [27].
@@
III Maximum harvesting rate
Suppose that the combined rate matrix $R=R^{b}+R^{c}$ has the steady-state distribution $\bm{\pi}$, which satisfies $R^{b}\bm{\pi}+R^{c}\bm{\pi}=0$. The total steady-state harvesting rate due to baseline and control is
@@ III Maximum harvesting rate
We seek to maximize this harvesting rate by varying the parameters of the control processes $(R^{c},\bm{g}^{c})$ while holding the baseline parameters ($\bm{f},R^{b},\bm{g}^{b},\dot{\bm{g}}^{b}$) fixed. Finding this maximum would allow us to determine fundamental bounds on harvesting and to evaluate the efficiency of existing harvesting systems.
However, $\dot{\mathcal{G}}^{\mathrm{tot}}$ is not a concave function of the parameters $(R^{c},\bm{g}^{c})$, therefore maximization of (7) is not a convex optimization problem and is not generally intractable. In the following, we reformulate this maximization as a convex optimization with a physically interpretable objective. This allows us to solve the optimization numerically and, at least for some special cases, also in closed form.
To begin, we rewrite (7) as
@@
where we introduced the Schnackenberg formula for the entropy production rate (EPR) [22],
$\dot{\Sigma}({\boldsymbol{J}}^{c})=\sum_{i\neq j}{J^{c}_{ji}}\ln({J^{c}_{ji}}/{J^{c}_{ij}})\geq 0\,,$ (9)
where ${J^{c}_{ji}}=\pi_{i}R^{c}_{ji}\geq 0$ is the one-way probability flux due to control transition $i\to j$.
Eq. (8) has an intuitive physical interpretation: the total steady-state harvesting rate is the rate of free energy increase in the system and internal reservoir due to baseline, minus the rate of dissipation (EPR) due to the control fluxes. The derivation of this result proceeds in two steps (see SM-I.1 [27] for details). The first step is to show that $\dot{\Sigma}({\boldsymbol{J}}^{c})=-\beta[\dot{\mathcal{F}}^{c}(\bm{\pi})+\beta\dot{\mathcal{G}}^{c}(\bm{\pi})]$, which follows by combining (9) with (3) and (6). This states that the EPR due to control is proportional to the free energy loss in the system and internal reservoir due to control. The second step is to show that ˙$\dot{\mathcal{F}}^{b}(\bm{\pi})+\dot{\mathcal{F}}^{c}(\bm{\pi})=0$, which follows because the left side is the overall derivative of the nonequilibrium free energy $\mathcal{F}$, as defined in (1), therefore it must vanish in steady state. The result (8) then follows by combining with (7) and rearranging.
Importantly, when expressed in the form (8), the harvesting rate is a concave function of the steady-state distribution $\bm{\pi}$ and the control fluxes ${\boldsymbol{J}}^{c}$ (see SM-I.2 [27]). To find the maximum harvesting rate, we optimize (8) with respect to $\bm{\pi}$ and ${\boldsymbol{J}}^{c}$. Note that varying $\bm{\pi}$ and ${\boldsymbol{J}}^{c}$ is equivalent to varying the control rate matrix via $R^{c}_{ji}={J^{c}_{ji}}/\pi_{i}$ and control driving ${g}^{c}_{ji}$ via (6). However, when performing this optimization, we must also ensure that $\bm{\pi}$ is the steady-state distribution induced by the fluxes ${\boldsymbol{J}}^{c}$. This condition can be expressed as a linear constraint on $\bm{\pi}$ and ${\boldsymbol{J}}^{c}$ via $R^{b}\bm{\pi}+\mathbb{B}{\boldsymbol{J}}^{c}=0$. Here ${\boldsymbol{J}}^{c}$ is treated as a vector in $\mathbb{R}^{n^{2}}$ and $\mathbb{B}\in\mathbb{R}^{n\times n^{2}}$ is the incidence matrix with entries $\mathbb{B}_{k,ij}=\delta_{ki}-\delta_{kj}$, which guarantees $R^{c}\bm{\pi}=\mathbb{B}{\boldsymbol{J}}^{c}$.
@@
$\mathscr{G}$ $=\sup_{(\bm{p},{\boldsymbol{J}})\in\Lambda:R^{b}\bm{p}+\mathbb{B}{\boldsymbol{J}}=0}\,\dot{\mathcal{F}}^{b}(\bm{p})+\dot{\mathcal{G}}^{b}(\bm{p})-\dot{\Sigma}({\boldsymbol{J}})/\beta.$ (10)
In this expression, $\Lambda$ is the feasible set of distributions $\bm{p}$ and control fluxes ${\boldsymbol{J}}$. At a minimum, $\Lambda$ ensures the validity of the distribution $\bm{p}$ and the fluxes ${\boldsymbol{J}}$ via the linear constraints $\sum p_{i}=1$, $p_{i}\geq 0$, and $J_{ji}\geq 0$. We write $\sup$ instead of $\max$ because the set of allowed fluxes is potentially unbounded. Eq. (10) implies a tradeoff between the total gain of free energy in the system and internal reservoir due to baseline (which depends only on $\bm{p}$) and the dissipation due to control fluxes (which depends only on ${\boldsymbol{J}}$).
Importantly, the feasible set $\Lambda$ can include additional convex constraints, which may introduce topological, kinetic, thermodynamic, etc. restrictions on the control processes. Topological constraints restrict which transitions can be controlled; e.g., $J_{ji}=0$ ensures that control does not use transition $i\to j$). Kinetic constraints restrict timescales of control processes, as might reflect underlying physical processes like diffusion; e.g., an upper bound on control transition rate $R^{c}_{ji}=J_{ji}/p_{i}\leq\kappa$ can be enforced via $J_{ji}\leq p_{i}\kappa$. Thermodynamic constraints bound the affinity [22] of control transitions; e.g., $J_{ji}e^{-\mathcal{A}}\leq J_{ij}\leq J_{ji}e^{\mathcal{A}}$ ensures that $|\ln(J_{ij}/J_{ji})|\leq\mathcal{A}$. The above examples all involve linear constraints. An example of a nonlinear, but still convex, constraint is an upper bound on the EPR incurred by control, $\dot{\Sigma}({\boldsymbol{J}})\leq\dot{\Sigma}^{c}_{\max}$.
Eq. (10) is our first main result. Importantly, $\mathscr{G}$ is defined purely in terms of the thermodynamic and kinetic properties of the baseline process, along with desired constraints encoded in $\Lambda$. Thus, $\mathscr{G}$ is the maximum steady-state harvesting rate that can be achieved by any allowed control processes, given a fixed baseline. In addition, Eq. (10) involves the maximization of a concave objective given convex constraints. This is equivalent to the minimization of a convex objective, thus Eq. (10) is a convex optimization problem that can be efficiently solved using standard numerical techniques [30]. The optimization also identifies an optimal steady-state distribution $\bm{p}^{*}$ and control fluxes ${{\boldsymbol{J}}}^{*}$ that achieve the maximum harvesting rate $\mathscr{G}$ (or come arbitrarily close to achieving it). These fix the optimal control rate matrix via $R^{c*}_{ji}={J_{ij}^{*}}/p^{*}_{i}$. Thus, our optimization specifies an upper bound on harvesting as well as the optimal control strategy that achieves this bound.
@@
The optimal $\bm{p}^{*}$ is unique as long as the baseline rate matrix is irreducible. The optimal control rate matrix is very fast ($R^{c*}\to\infty$) and obeys detailed balance for $\bm{p}^{*}$, $R^{c*}_{ji}{p}^{*}_{i}=R^{c*}_{ij}{p}^{*}_{j}$. For details, see SM-I.3 and SM-I.4 [27].
IV Bacteriorhodopsin
We illustrate our results using bacteriorhodopsin, a light-driven proton pump from Archaea [18].
We define a thermodynamically consistent model of the bacteriorhodopsin cycle using published thermodynamic [31] and kinetic [32] data (see SM-II [27]). The system operates in a cyclical manner, absorbing a photon and pumping a proton during each turn of the cycle (Fig. 1, right). Specifically, the transition $M_{1}\to M_{2}$ pumps a proton out of the cell. This stores free energy in the internal reservoir (the membrane electrochemical potential),